OpenStudy (anonymous):

Solve for x: 3|2x − 2| + 6 = 18. x = 3, x = −3 x = 3, x = −1 x = 1, x = −1 x = −3, x = 1

OpenStudy (anonymous):

@imqwerty

pooja195 (pooja195):

Add -6 to both sides. 3(|2x−2|)+6+−6=18+−6 3(|2x−2|)=12 divide both sides by 3 we are left with this |2x−2|=4

pooja195 (pooja195):

We know either 2x−2=4 or 2x−2=−4

pooja195 (pooja195):

2x-2=4 solve this for x

OpenStudy (anonymous):

x=3 ?

pooja195 (pooja195):

Yes so one of out answer is x=3 now solve this for x 2x-2=-4

OpenStudy (anonymous):

answer is x=3 , x=-1 ?

pooja195 (pooja195):

Yes :) well done

OpenStudy (anonymous):

Thank You

pooja195 (pooja195):

No problem